Let a(s) = (x(s), y(s), 0) be a unit speed curve. Prove that


K = |x'y" – x''y'l.


1
Expert's answer
2021-04-12T15:44:00-0400

Consider the identity

"(\dot{x}^2+\dot{y}^2)(\ddot{x}^2+\ddot{y}^2)=(\dot{x}\ddot{x}+\dot{y}\ddot{y})^2+(\dot{x}\ddot{y}-\ddot{x}\dot{y})^2"

With x=x(s), y=y(s) and dots denoting differentiation over the variable s, we have

"\dot{x}^2+\dot{y}^2=1" by condition

"\ddot{x}^2+\ddot{y}^2=|\ddot{a}(s)|^2=K^2" by definition

"\dot{x}\ddot{x}+\dot{y}\ddot{y}=\frac{1}{2}\frac{d}{ds}(\dot{x}^2+\dot{y}^2)=\frac{1}{2}\frac{d}{ds}(1)=0"

Then we obtain "K^2=(\dot{x}\ddot{y}-\ddot{x}\dot{y})^2" or, equivalently, "K=|\dot{x}\ddot{y}-\ddot{x}\dot{y}|", as required.